# Conformalized Polynomial Chaos Expansion for Uncertainty-aware Surrogate Modeling

### v2

@@ 2.1 Regression-based polynomial chaos expansion
Next, we introduce a multivariate random variable (or random vector) $\mathbf{X}=\left(X_{1},\dots,X_{N}\right)$, where the individual random variables $X_{n}$, $n=1,\dots,N$, are assumed to be mutually independent. The random vector is defined on the probability space $\left(\Omega,\Sigma,P\right)$, where $\Omega$ is the sample space containing all measurable outcomes $\omega\in\Omega$, $\Sigma$ denotes the $\sigma$-algebra of events (grouped outcomes), and $P:\Sigma\rightarrow[0,1]$ is a probability measure. The latter is defined based on the assumed probability density function (PDF) $f_{\mathbf{X}}:\Xi\rightarrow\mathbb{R}_{\geq 0}$, where $\Xi\subset\mathbb{R}^{N}$ denotes the state-space of the random vector. Then, the random vector is a map, $\mathbf{X}:\Omega\rightarrow\Xi$, with random realizations $\mathbf{x}=\mathbf{X}\left(\omega\right)\in\Xi$.
@@
Given a multi-index set $\Lambda$ and the corresponding polynomial basis, the remaining task is to estimate the PCE coefficients $\mathbf{c}=\left(c_{\boldsymbol{\alpha}}\right)_{\boldsymbol{\alpha}\in\Lambda}^{\top}\in\mathbb{R}^{K\times 1}$. We opt here for a regression-based approach [migliorati2013approximation, migliorati2014analysis, hadigol2018least]. Alternatives include include pseudo-spectral projection [constantine2012sparse, conrad2013adaptive] or interpolation [buzzard2012global, porta2014inverse, galetzka2023hp]. For the regression-based coefficient estimation, we assume a set of input parameter realizations along with the corresponding model responses $\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$, commonly called the experimental design or, in an ML-context, the training data set. We then construct the so-called design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, such that $d_{mk}=\Psi_{k}\left(\mathbf{x}^{(m)}\right)$. Collecting the model responses in a vector $\mathbf{y}=\left(y^{(1)},\dots,y^{(M)}\right)^{\top}\in\mathbb{R}^{M\times 1}$, the coefficients can be computed by solving the least squares minimization problem
$\mathbf{c}=\underset{\hat{\mathbf{c}}\in\mathbb{R}^{K}}{\arg\min}\left\{\frac{1}{M}\lVert\mathbf{y}-\mathbf{D}\hat{\mathbf{c}}\rVert_{2}\right\},$ (4)
@@ 2.2 Conformal prediction
Following the notation introduced in section 2.1, we consider a regression model $\widehat{\mu}\left(\mathbf{x}\right)$ trained with the dataset $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. Then, given a test point $\left\{\mathbf{x}^{*},y^{*}\right\}\not\in\mathcal{D}_{\text{train}}$, the trained model produces a prediction $\widehat{y}^{*}=\widehat{\mu}(\mathbf{x}^{*})$ for the true value $y^{*}$. The goal of CP is to complement the model’s prediction with an interval which is likely to contain the true output $y^{*}$. More formally, given a significance level $s\in\left(0,1\right)$ corresponding to the target coverage level (1s)(1-s), CP constructs a prediction interval $\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$, such that
@@
such that $y^{*}\in\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$ implies that $\alpha\!\left(\mathbf{x}^{*},y^{*};\widehat{\mu}\right)\leq\widehat{q}_{1-s}\!\left\{\alpha_{j}\right\}$ and vice versa. A commonly used non-conformity score is the absolute residual, such that $\alpha_{j}=\left|r_{j}\right|=\left|y^{(j)}-\widehat{\mu}\left(\mathbf{x}^{(j)}\right)\right|$. This choice results in the construction of the symmetric predictive interval
$\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)=\left[\widehat{\mu}\left(\mathbf{x}^{*}\right)-\widehat{q}_{1-s}\left\{\left|r_{j}\right|\right\},\widehat{\mu}\left(\mathbf{x}^{*}\right)+\widehat{q}_{1-s}\left\{\left|r_{j}\right|\right\}\right].$ (8)
@@
A naive approach to estimate the predictive interval (8) would be to use the residuals on the training data $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. However, training data residuals are expected to be smaller compared to residuals obtained from previously unseen test data points. Hence, the resulting predictive interval will most probably not provide the required coverage. Split or inductive CP approaches [papadopoulos2002inductive, lei2018distribution] use a separate validation dataset $\mathcal{D}_{\text{val}}\neq\mathcal{D}_{\text{train}}$ to compute the non-conformity scores and construct the prediction interval (8). Split CP guarantees the required finite-sample coverage (6), however, this comes at the cost of a larger data demand. To avoid the need for additional data, cross CP [vovk2015cross, vovk2018cross] utilizes $K$-fold cross-validation, such that the training dataset $\mathcal{D}_{\text{train}}$ is split into $K$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{K}$ of (approximately) equal size. Then, the regression models $\widehat{\mu}_{\sim k}\left(\mathbf{x}\right)$ are trained using the datasets $\mathcal{D}_{\sim k}=\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{k}$, $k=1,\dots,K$, and the cross-validation residuals
$r_{m}^{\text{cv}}=y^{(m)}-\widehat{\mu}_{\sim k(m)}\left(\mathbf{x}^{(m)}\right),$ (9)
are used as non-conformity scores (in absolute value), where the index $k(m)$ denotes that $\left(\mathbf{x}^{(m)},y^{(m)}\right)\in\mathcal{D}_{k}$. Cross CP typically yields an empirical coverage close to the target level (1s)(1-s), but does not provide formal coverage guarantees. A major downside of cross CP is that $K$ different regression models must be trained, which can lead to an undesired computational cost.
A special case of cross-validation is the jackknife method [chernick2012jackknife]. Using the jackknife in the context of CP [lei2018distribution], the non-conformity scores coincide with the absolute values of the LOO residuals
$r_{m}^{\text{LOO}}=y^{(m)}-\widehat{\mu}_{\sim m}\left(\mathbf{x}^{(m)}\right),$ (10)
@@
$\widehat{C}_{1-s}^{\text{jack}}\left(\mathbf{x}^{*}\right)$ $=\left[\widehat{\mu}\left(\mathbf{x}^{*}\right)-\widehat{q}_{1-s}\left\{\left|r_{m}^{\text{LOO}}\right|\right\},\widehat{\mu}\left(\mathbf{x}^{*}\right)+\widehat{q}_{1-s}\left\{\left|r_{m}^{\text{LOO}}\right|\right\}\right]$ (11a)
$=\left[\widehat{q}_{s}\left\{\widehat{\mu}\left(\mathbf{x}^{*}\right)-\left|r_{m}^{\text{LOO}}\right|\right\},\widehat{q}_{1-s}\left\{\widehat{\mu}\left(\mathbf{x}^{*}\right)+\left|r_{m}^{\text{LOO}}\right|\right\}\right].$ (11b)
@@
$C_{1-s}^{\text{jack+}}\left(\mathbf{x}^{*}\right)=\left[\widehat{q}_{s}\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)-\left|r_{m}^{\text{LOO}}\right|\right\},\widehat{q}_{1-s}\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)+\left|r_{m}^{\text{LOO}}\right|\right\}\right].$ (12)
@@ 2.4 Normalized non-conformity scores
In the previous section, the non-conformity scores were equal to the absolute LOO residuals, i.e., $\alpha_{m}=\left|r_{m}^{\text{LOO}}\right|$, $m=1,\dots,M$. Based on the relative LOO error, defined as
@@
$\alpha_{m}^{\text{norm}}=\frac{\left|r_{m}^{\text{LOO}}\right|}{\sqrt{\text{Var(Y)}}}.$ (17)
@@
The main advantage of using normalized non-conformity scores is that they adapt to heteroskedastic data, thus offering prediction intervals that vary in width according to local uncertainty. The main downside is that an output variance estimate is required, which may be challenging. For a PCE model, the output variance can be estimated cost-effectively by post-processing its coefficients, such that
@@
assuming an orthonormal polynomial basis. Note that this estimate is affected by the accuracy of the PCE. Alternatively, the ouput variance can be estimated from the experimental design, assuming that there exist enough data points to allow for a sufficiently accurate estimate.
@@ 4.  Summary, conclusions, and outlook
The conformalized PCE method developed in this work shows definite promise as uncertainty-aware, data-driven, surrogate modeling method. Nonetheless, it can be improved in several aspects. First, this work has considered total-degree PCEs only, thus limiting the selected benchmark models and engineering use-cases to low dimensions and/or polynomial degrees. The integration of CP within sparse PCE algorithms [luethen2021sparse, luethen2022automatic] would allow to extend the method to more challenging problem settings. Second, in its current form, the conformalized PCE method is applicable to scalar outputs only. In principle, it can be applied elementwise to address multidimensional outputs. However, this would be a suboptimal approach, as it does not account for correlations among the output’s elements and would incur high computational costs with increasing output dimension. The development of a conformalized PCE method suitable for multivariate outputs seems to be a natural follow-up, for example based on the combination of multi-output PCEs [hawchar2017principal, bhattacharyya2020uncertainty, loukrezis2025multivariate] and multi-output CP [zhang2023improved, xu2023conformal]. We aim to address these limitations in dedicated follow-up works.

### v3

@@ 2.1 Regression-based polynomial chaos expansion
Next, we introduce a multivariate random variable (or random vector) $\mathbf{X}=\left(X_{1},\dots,X_{N}\right)$, where the individual random variables $X_{n}$, $n=1,\dots,N$, are assumed to be mutually independent. The random vector is defined on the probability space $\left(\Omega,\Sigma,P\right)$, where $\Omega$ is the sample space containing the outcomes $\omega\in\Omega$, $\Sigma$ denotes the $\sigma$-algebra of events (measurable grouped outcomes), and $P:\Sigma\rightarrow[0,1]$ is a probability measure. The latter is defined based on the assumed probability density function (PDF) $f_{\mathbf{X}}:\Xi\rightarrow\mathbb{R}_{\geq 0}$, where $\Xi\subset\mathbb{R}^{N}$ denotes the state-space of the random vector. Then, the random vector is a map, $\mathbf{X}:\Omega\rightarrow\Xi$, with random realizations $\mathbf{x}=\mathbf{X}\left(\omega\right)\in\Xi$.
@@
Given a multi-index set $\Lambda$ and the corresponding polynomial basis, the remaining task is to estimate the PCE coefficients $\mathbf{c}=\left(c_{\boldsymbol{\alpha}}\right)_{\boldsymbol{\alpha}\in\Lambda}^{\top}\in\mathbb{R}^{K\times 1}$. We opt here for a regression-based approach [migliorati2013approximation, migliorati2014analysis, hadigol2018least]. Alternatives include include pseudo-spectral projection [constantine2012sparse, conrad2013adaptive] or interpolation [buzzard2012global, porta2014inverse, galetzka2023hp]. For the regression-based coefficient estimation, we assume a set of input parameter realizations $\left\{\mathbf{x}^{(m)}\right\}_{m=1}^{M}$, commonly called the experimental design, along with the corresponding model responses $\left\{y^{(m)}\right\}_{m=1}^{M}$. Put in an ML context, the combined dataset $\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$ constitutes the training dataset. We then construct the so-called design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, such that $d_{mk}=\Psi_{k}\left(\mathbf{x}^{(m)}\right)$. Collecting the model responses in a vector $\mathbf{y}=\left(y^{(1)},\dots,y^{(M)}\right)^{\top}\in\mathbb{R}^{M\times 1}$, the coefficients can be computed by solving the least squares minimization problem
$\mathbf{c}=\underset{\hat{\mathbf{c}}\in\mathbb{R}^{K}}{\arg\min}\left\{\frac{1}{M}\lVert\mathbf{y}-\mathbf{D}\hat{\mathbf{c}}\rVert_{2}^{2}\right\},$ (4)
@@ 2.2 Conformal prediction
Following the notation introduced in section 2.1, we consider a regression model $\widehat{\mu}\left(\mathbf{x}\right)$ trained with the dataset $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$, where again the experimental design data points $\mathbf{x}^{(m)}$ are assumed to be realizations of random vector $\mathbf{X}$ characterized by a joint probability distribution $f_{\mathbf{X}}(\mathbf{x})$. Then, given a test point $\left\{\mathbf{x}^{*},y^{*}\right\}\not\in\mathcal{D}_{\text{train}}$, the trained model produces a prediction $\widehat{y}^{*}=\widehat{\mu}(\mathbf{x}^{*})$ for the true value $y^{*}$. The goal of CP is to complement the model’s prediction with an interval which is likely to contain the true output $y^{*}$. More formally, given a significance level $s\in\left(0,1\right)$ corresponding to the target coverage level (1s)(1-s), CP constructs a prediction interval $\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$, such that
@@
such that $y^{*}\in\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$ implies that $\alpha\!\left(\mathbf{x}^{*},y^{*};\widehat{\mu}\right)\leq\widehat{q}_{1-s}\!\left\{\alpha_{j}\right\}$ and vice versa. A commonly used non-conformity score is the absolute residual, such that $\alpha_{j}=R_{j}=\left|r_{j}\right|=\left|y^{(j)}-\widehat{\mu}\left(\mathbf{x}^{(j)}\right)\right|$. This choice results in the construction of the symmetric predictive interval
$\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)=\left[\widehat{\mu}\left(\mathbf{x}^{*}\right)-\widehat{q}_{1-s}\left\{R_{j}\right\},\widehat{\mu}\left(\mathbf{x}^{*}\right)+\widehat{q}_{1-s}\left\{R_{j}\right\}\right].$ (8)
@@
A naive approach to estimate the predictive interval (8) would be to use the residuals on the training data $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. However, training data residuals are expected to be smaller compared to residuals obtained from previously unseen test data points. Hence, the resulting predictive interval will most probably not provide the required coverage. Split or inductive CP approaches [papadopoulos2002inductive, lei2018distribution] use a separate validation dataset $\mathcal{D}_{\text{val}}\neq\mathcal{D}_{\text{train}}$ to compute the non-conformity scores and construct the prediction interval (8). Split CP guarantees the required finite-sample coverage (6), however, this comes at the cost of a larger data demand. To avoid the need for additional data, cross CP [vovk2015cross, vovk2018cross] utilizes $K$-fold cross-validation, such that the training dataset $\mathcal{D}_{\text{train}}$ is split into $K$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{K}$ of (approximately) equal size. Then, the regression models $\widehat{\mu}_{\sim k}\left(\mathbf{x}\right)$ are trained using the datasets $\mathcal{D}_{\sim k}=\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{k}$, $k=1,\dots,K$, and the absolute cross-validation residuals
$R_{m}^{\text{cv}}=\left|r_{m}^{\text{cv}}\right|=\left|y^{(m)}-\widehat{\mu}_{\sim k(m)}\left(\mathbf{x}^{(m)}\right)\right|,$ (9)
are used as non-conformity scores, where the index $k(m)$ denotes that $\left(\mathbf{x}^{(m)},y^{(m)}\right)\in\mathcal{D}_{k}$. Cross CP typically yields an empirical coverage close to the target level (1s)(1-s), but does not provide formal coverage guarantees. A major downside of cross CP is that $K$ different regression models must be trained, which can lead to an undesired computational cost.
A special case of cross-validation is the jackknife method [chernick2012jackknife]. Using the jackknife in the context of CP [lei2018distribution], the non-conformity scores coincide with the absolute LOO residuals
$R_{m}^{\text{LOO}}=\left|r_{m}^{\text{LOO}}\right|=\left|y^{(m)}-\widehat{\mu}_{\sim m}\left(\mathbf{x}^{(m)}\right)\right|,$ (10)
@@
$\widehat{C}_{1-s}^{\text{jack}}\left(\mathbf{x}^{*}\right)$ $=\left[\widehat{\mu}\left(\mathbf{x}^{*}\right)-\widehat{q}_{1-s}\left\{R_{m}^{\text{LOO}}\right\},\widehat{\mu}\left(\mathbf{x}^{*}\right)+\widehat{q}_{1-s}\left\{R_{m}^{\text{LOO}}\right\}\right]$ (11a)
$=\left[\widehat{q}_{s}\left\{\widehat{\mu}\left(\mathbf{x}^{*}\right)-R_{m}^{\text{LOO}}\right\},\widehat{q}_{1-s}\left\{\widehat{\mu}\left(\mathbf{x}^{*}\right)+R_{m}^{\text{LOO}}\right\}\right].$ (11b)
@@
$\widehat{C}_{1-s}^{\text{jack+}}\left(\mathbf{x}^{*}\right)=\left[\widehat{q}_{s}\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)-R_{m}^{\text{LOO}}\right\},\widehat{q}_{1-s}\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)+R_{m}^{\text{LOO}}\right\}\right].$ (12)
@@ 2.4 Normalized non-conformity scores
In the previous section, the non-conformity scores were equal to the absolute LOO residuals, i.e., $\alpha_{m}=R_{m}^{\text{LOO}}=\left|r_{m}^{\text{LOO}}\right|$, $m=1,\dots,M$. Based on the relative LOO error, defined as
@@
$\alpha_{m}^{\text{norm}}=\frac{R_{m}^{\text{LOO}}}{\sqrt{\text{Var(Y)}}}.$ (17)
@@
The main advantage of using these normalized non-conformity scores is that they are scale-invariant, thus possibly resulting in more stable quantile estimates. The main downside is that an output variance estimate is required. For a PCE model, the output variance can be estimated cost-effectively by post-processing its coefficients, such that
@@
assuming an orthonormal polynomial basis. However, this estimate is affected by the accuracy of the PCE. Alternatively, the ouput variance can be estimated from the experimental design, assuming that there exist enough data points to allow for a sufficiently accurate estimate.
Note that the normalized non-conformity scores given by formula (17) still result in symmetric predictive intervals. Therefore, they should not be confused with non-conformity scores normalized by the conditional mean absolute deviation of the model’s prediction, also called “normalized” in the literature [lei2018distribution], which result in locally weighted intervals that adapt to heteroskedastic data.
@@ 4.  Summary, conclusions, and outlook
The conformalized PCE method developed in this work shows definite promise as uncertainty-aware, data-driven, surrogate modeling method. Nonetheless, it can be improved in several aspects. First, this work has considered total-degree PCEs only, thus limiting the selected benchmark models and engineering use-cases to low dimensions and/or polynomial degrees. The integration of CP within sparse PCE algorithms [luethen2021sparse, luethen2022automatic] would allow to extend the method to more challenging problem settings. Second, in its current form, the conformalized PCE method is applicable to scalar outputs only. In principle, it can be applied elementwise to address multidimensional outputs. However, this would be a suboptimal approach, as it does not account for correlations among the output’s elements and would incur high computational costs with increasing output dimension. The development of a conformalized PCE method suitable for multivariate outputs seems to be a natural follow-up, for example based on the combination of multi-output PCEs [hawchar2017principal, bhattacharyya2020uncertainty, loukrezis2025multivariate] and multi-output CP [zhang2023improved, xu2023conformal]. Third, the current approach provides symmetric predictive intervals that do not adapt to heteroskedastic data. Combinations with locally weighted non-conformity measures [lei2018distribution] or with conformalized quantile regression [romano2019conformalized] could offer possible remedies. We aim to address these limitations in dedicated follow-up works.